OpenStudy (anonymous):

describe the end behavior 2x5-7x2-4x

OpenStudy (anonymous):

\[f(x)=2x^5-7x^2-4x\] To describe the end behavior you want to see what happens as x -> infinity and -infinity. \[\lim_{x \rightarrow \infty}f(x) \rightarrow \infty \] \[\lim_{x \rightarrow -\infty}f(x)\rightarrow-\infty \] So the function goes to infinity towards the right and goes towards negative infinity towards the left.

OpenStudy (anonymous):

use the synthetic division to evaluate f(x)=z3-8z+6 when z=-3